OpenStudy (hba):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (hba):

@hartnn @lgbasallote @Compassionate

hartnn (hartnn):

asked here so many times...i'll let others try.....

OpenStudy (hba):

@klimenkov

OpenStudy (lgbasallote):

i'll let @hartnn take this one. im interested to see if he really knows

hartnn (hartnn):

this is pure algebra...2 equations 2 unknowns.........

OpenStudy (hba):

Do it

OpenStudy (lgbasallote):

algebra is boring. it's much more fun using common sense and logical reasoning

OpenStudy (theviper):

didn't tag me :\

OpenStudy (hba):

@TheViper

OpenStudy (hba):

tagged

OpenStudy (theviper):

LOL

OpenStudy (hba):

@UsukiDoll

OpenStudy (compassionate):

First start off by making a system of linear equations: x + y = z 2x + 4y = 196

OpenStudy (compassionate):

x + y = 74

hartnn (hartnn):

x+y=74 2x+4y=196

OpenStudy (klimenkov):

horses=24 people=50

OpenStudy (compassionate):

^ Correct. Well, question answered.

OpenStudy (hba):

How did u get these equations

hartnn (hartnn):

50 and 24 indeed

OpenStudy (lgbasallote):

you can just "calculate" this way 74 heads so 74 one legs.. then if you add 74 more 148 two legs so there are at most 148 humans. howerever, 196 - 148 = 48 so there are 48 more legs. so if you subtract 48 from 148 you get 100 legs <--humans so divide by 2 so there are 50 men

OpenStudy (lgbasallote):

logical reasoning is much more fun than letters

OpenStudy (hba):

Thanks ALot Lgba :)

OpenStudy (compassionate):

Well, I was going to solve it algebraically.

OpenStudy (lgbasallote):

pleasures of hating math

hartnn (hartnn):

the difference between algebraic and logical is: algebra use x,y,sum,difference logical only use sum,difference.........

OpenStudy (hba):

Does Anyone in here know about Fermat's Last Theorem ?

OpenStudy (lgbasallote):

logic requires thinking and manipulation algebra makes logic easier by using representations

hartnn (hartnn):

then i would go for algebra....

OpenStudy (lgbasallote):

of course

OpenStudy (lgbasallote):

you're a math lover after all

OpenStudy (hba):

Haha LoL

hartnn (hartnn):

*proudly* YES :D